# Efficient Testing Implies Structured Symmetry

### v1

@@ 2 Preliminaries
A property $\mathcal{P}$ is a set of Boolean functions $f:\{0,1\}^{n}\to\{0,1\}$. We say that $f$ has the property $\mathcal{P}$ if $f\in\mathcal{P}$. The distance between two functions is the fraction of inputs on which they disagree. We write $f\in\mathcal{P}_{\varepsilon}$ if $f$ is $\varepsilon$-close to some $g\in\mathcal{P}$. While property testing can be studied in either a query-based or sample-based access model, we focus exclusively on the latter perspective, which was introduced by [GGR96]:
@@
A property $\mathcal{P}$ of Boolean functions is $k$-part symmetric if there is a partition of $\{0,1\}^{n}$ into disjoint parts $S_{1},\ldots,S_{k}$ such that $\mathcal{P}$ is invariant under permutations of the domain within each part. Equivalently, whether or not a function $f$ has the property $\mathcal{P}$ can be completely determined from the $k$ scalar densities $\operatorname*{\mathbb{E}}[f(x)\,|\,x\in S_{j}]$, where the expectation is computed over a random input $x\sim\{0,1\}^{n}$. In general, such parts need not have any special structure, and may be very complex. In contrast, we say that a part $S_{j}$ has size complexity at most $s$ if there is a circuit of size $s$ that decides whether or not a given input belongs to $S_{j}$. We say that a partition $\mathcal{P}$ has size complexity at most $s$ if there is a circuit of size at most $s$ computing its classification function, which computes the index of the part $S_{j}$ containing a given input $x$.
Supersimulators
Our key technical tool is the construction of supersimulators [DT25], which strengthens the complexity-theoretic regularity lemma [TTV09] by designing regular simulators that fool distinguishers more powerful than themselves. First, we define indistinguishability formally.
@@
Given a family $\mathcal{F}\subseteq\{\mathcal{X}\to[0,1]\}$, an error parameter $\delta>0$, and two functions $g,h:\mathcal{X}\to[0,1]$, we say that $h$ is a $(\mathcal{F},\delta)$-regular simulator for $g$ under $\mathcal{D}$ if for all distinguisher functions $f\in\mathcal{F}$,
@@
If $\mathcal{F}\subseteq\{\mathcal{X}\to[-1,1]\}$, then $\mathcal{S}_{k,\delta}(\mathcal{F})$ is the set of functions
$h(x)=\Bigl[\delta\cdot\bigl(f_{1}(x)+\cdots+f_{k}(x)\bigr)\Bigr]_{0}^{1}$
for some $f_{1},\ldots,f_{k}\in\pm\mathcal{F}$. (Recall that $\cdot]_{0}^{1}$ projects onto $[0,1]$.) Let $\mathcal{S}_{<k,\delta}(\mathcal{F})=\bigcup_{j<k}\mathcal{S}_{j,\delta}(\mathcal{F})$.
@@
Fix $\mathcal{D}\in\Delta(\mathcal{X})$, $\mathcal{F}\subseteq\{\mathcal{X}\to[-1,1]\}$, and $\delta>0$. Every $g:\mathcal{X}\to[0,1]$ has an $(\mathcal{F},\delta)$-regular simulator $h\in\mathcal{S}_{<(2/\delta^{2}),(\delta/2)}(\mathcal{F})$.
@@
Lemma 2.4 may produce a simulator $h$ that is more complex than the functions in $\mathcal{F}$, but for our purposes, we will need $h$ to be a supersimulator, which fools distinguishers more complex than itself. We formalize this in terms of a growth function $\mathcal{G}$ that takes as input any function $h:\mathcal{X}\to[0,1]$ and outputs the distinguisher family $\mathcal{G}(h)\subseteq\{\mathcal{X}\to[-1,1]\}$ that we would like $h$ to fool. Now, the definition of structured sums must be adjusted accordingly.
Definition 2.5.
If $\mathcal{G}$ is a growth function, then $\mathcal{S}_{k,\delta}(\mathcal{G})$ is the set of functions
$h_{k}(x)=\Bigl[\delta\cdot\bigl(f_{1}(x)+\cdots+f_{k}(x)\bigr)\Bigr]_{0}^{1}$
@@
where each $f_{j}$ belongs to the previous family $\pm\mathcal{G}(h_{j-1})$, and $h_{0}=0$. Let $\mathcal{S}_{<k,\delta}(\mathcal{G})=\bigcup_{j<k}\mathcal{S}_{j,\delta}(\mathcal{G})$.
Lemma 2.6 (Supersimulators [DT25]).
@@
Fix $\mathcal{D}\in\Delta(\mathcal{X})$, a growth function $\mathcal{G}$, and $\delta>0$. Every $g:\mathcal{X}\to[0,1]$ has a $(\mathcal{G}(h),\delta)$-regular simulator $h\in\mathcal{S}_{<(2/\delta^{2}),(\delta/2)}(\mathcal{G})$.
Note that Lemma 2.4 is a special case of Lemma 2.6 corresponding to a constant growth function that always outputs $\mathcal{F}$. We remark that the statement of Lemma 2.6 differs slightly from that of [DT25] because Lemma 2.6 uses a single projection onto the interval $[0,1]$, as opposed to a nested sequence of projections. For completeness, we provide a short proof of this version in Appendix A. In fact, our argument also yields a slightly simpler proof of Lemma 2.4 than in [TTV09], which required some careful case analysis to handle the projection. Simple proofs of the complexity-theoretic regularity lemma without this case analysis have already appeared (e.g. implicitly in [HKRR18], as observed by [CDV24]), but these versions also require a sequence of nested projections, unlike Lemma 2.4, which more closely resembles the original version from [TTV09].

### v2

@@ 2 Preliminaries
A property $\mathcal{P}$ is a set of Boolean functions $f:\{0,1\}^{n}\to\{0,1\}$. We say that $f$ has the property $\mathcal{P}$ if $f\in\mathcal{P}$. The distance between two functions is the fraction of inputs on which they disagree. We write $f\in\mathcal{P}_{\varepsilon}$ if $f$ is $\varepsilon$-close to some $g\in\mathcal{P}$. While property testing can be studied in either a query-based or sample-based access model, we focus on the latter perspective, which was introduced by [GGR96]:
@@
A property $\mathcal{P}$ of Boolean functions is $k$-part symmetric if there is a partition of $\{0,1\}^{n}$ into disjoint parts $S_{1},\ldots,S_{k}$ such that $\mathcal{P}$ is invariant under permutations of the domain within each part. Equivalently, whether or not a function $f$ has the property $\mathcal{P}$ can be completely determined from the $k$ scalar densities $\operatorname*{\mathbb{E}}[f(x)\bm{1}[x\in S_{j}]]$, where the expectation is computed over a random input $x\sim\{0,1\}^{n}$.
In general, such parts need not have any special structure, and may be very complex. In contrast, we say that a part $S_{j}$ has computational complexity at most $s$ if there is a circuit of size $s$ that decides whether or not a given input belongs to $S_{j}$. We say that a partition $\mathcal{P}$ has computational complexity at most $s$ if there is a circuit of size at most $s$ computing its classification function, which outputs the index of the unique part $S_{j}$ containing a given input $x$.
Complexity-Theoretic Regularity
First, we define regularity and indistinguishability.
@@
Given a family $\mathcal{F}\subseteq\{\mathcal{X}\to[-1,1]\}$, an error parameter $\delta>0$, and two functions $g,h:\mathcal{X}\to[0,1]$, we say that $h$ is a $(\mathcal{F},\delta)$-regular simulator for $g$ under $\mathcal{D}$ if for all distinguisher functions $f\in\mathcal{F}$,
@@
Given a distinguisher family $\mathcal{F}\subseteq\{\mathcal{X}\to[-1,1]\}$ and an initialization function $h_{0}:\mathcal{X}\to[0,1]$, the set $\mathcal{S}_{k,\delta}(\mathcal{F},h_{0})$ comprises all functions of the form
$h(x)=\Bigl[h_{0}(x)+\delta\cdot\bigl(f_{1}(x)+\cdots+f_{k}(x)\bigr)\Bigr]_{0}^{1}$
for some $f_{1},\ldots,f_{k}\in\pm\mathcal{F}$. [^10] Recall that $\cdot]_{0}^{1}$ projects onto $[0,1]$. Let $\mathcal{S}_{<k,\delta}(\mathcal{F};h_{0})=\bigcup_{j<k}\mathcal{S}_{j,\delta}(\mathcal{F},h_{0})$.
We will often use the initialization $h_{0}=0$. When this is clear from context, we will simply write $\mathcal{S}_{k,\delta}(\mathcal{F})$ and $\mathcal{S}_{<k,\delta}(\mathcal{F})$ in place of $\mathcal{S}_{k,\delta}(\mathcal{F},0)$ and $\mathcal{S}_{<k,\delta}(\mathcal{F},0)$.
@@
Fix $\mathcal{D}\in\Delta(\mathcal{X})$, $\mathcal{F}\subseteq\{\mathcal{X}\to[-1,1]\}$, $h_{0}:\mathcal{X}\to[0,1]$ and $\delta>0$. Every $g:\mathcal{X}\to[0,1]$ has an $(\mathcal{F},\delta)$-regular simulator $h\in\mathcal{S}_{<(2/\delta^{2}),(\delta/2)}(\mathcal{F},h_{0})$.
In the introduction, we had focused on the special case where $\mathcal{D}$ is the uniform distribution on $\mathcal{X}=\{0,1\}^{n}$ and $\mathcal{F}$ contains all circuits of size at most $s$, but this need not be the case. We had also considered randomized distinguishers $A(x,y)$ that take as a second input a binary label $y$ sampled from either $\mathcal{B}(g(x))$ or $\mathcal{B}(h(x))$, but here we consider deterministic distinguishers $f(x)$. Of course, the latter perspective subsumes the former. Given $A$, simply set $f_{A}(x)=\operatorname*{\mathbb{E}}[A(x,1)-A(x,0)]$, where the expectation is taken over the randomness of $A$. Then, the distinguishing advantage of $A$ is precisely $\lvert\operatorname*{\mathbb{E}}_{x\sim\mathcal{D}}[f_{A}(x)(g(x)-h(x))]\rvert$.
@@
The complexity-theoretic regularity lemma, as stated in Lemma 2.4, may produce a simulator $h$ that is more complex than the functions in $\mathcal{F}$, but for our purposes, we will need $h$ to be a supersimulator, which fools distinguishers more complex than itself. We formalize this in terms of a growth function $\mathcal{G}$ that takes as input any function $h:\mathcal{X}\to[0,1]$ and outputs the distinguisher family $\mathcal{G}(h)\subseteq\{\mathcal{X}\to[-1,1]\}$ that we would like $h$ to fool. Now, the notion of structured sums from Definition 2.3 must be adjusted accordingly.
Definition 2.5 (Structured Sums).
Fix $\delta>0$. Given an initialization $h_{0}:\mathcal{X}\to[0,1]$ and distinguisher functions $f_{1},f_{2},\ldots:\mathcal{X}\to[-1,1]$, let $f_{1:0}=h_{0}$ and for each $j\in[k]$, let
$f_{1:j}(x)=\Bigl[h_{0}(x)+\delta\cdot\bigl(f_{1}(x)+\cdots+f_{j}(x)\bigr)\Bigr]_{0}^{1}.$
@@
If $\mathcal{G}$ is a growth function, then $\mathcal{S}_{k,\delta}(\mathcal{G},h_{0})$ is the set of functions of the form of $f_{1:k}$ for some $f_{1},f_{2},\ldots$ such that each $f_{j}$ belongs to the previous family $\pm\mathcal{G}(f_{1:j-1})$. Let $\mathcal{S}_{<k,\delta}(\mathcal{G},h_{0})=\bigcup_{j<k}\mathcal{S}_{j,\delta}(\mathcal{G},h_{0})$.
Once again, we omit the second argument $h_{0}$ when it is clear from context that $h_{0}=0$.
Lemma 2.6 (Supersimulators).
@@
Fix $\mathcal{D}\in\Delta(\mathcal{X})$, a growth function $\mathcal{G}$, an initialization $h_{0}:\mathcal{X}\to[0,1]$, and $\delta>0$. Every $g:\mathcal{X}\to[0,1]$ has a $(\mathcal{G}(h),\delta)$-regular simulator $h\in\mathcal{S}_{<(2/\delta^{2}),(\delta/2)}(\mathcal{G},h_{0})$.
Note that Lemma 2.4 is a special case of Lemma 2.6 corresponding to a constant growth function that always outputs $\mathcal{F}$. While supersimulators appeared implicitly in [DKR+21], the statement of Lemma 2.6 differs slightly from the version in [DKR+21]. For this reason, we give a short proof of Lemma 2.6 in Appendix A, along with our alternate construction of supersimulators using an iteration technique from the graph regularity literature.
Multicalibration
A simulator (a.k.a. predictor) $h$ is said to be calibrated if $\operatorname*{\mathbb{E}}_{x\sim\mathcal{D}}[g(x)|h(x)]=h(x)$. Although the simulator provided by Lemma 2.4 is not necessarily perfectly calibrated, one can ensure that it is approximately calibrated at the cost of a small increase in the complexity of $h$. Specifically, following [GKSZ22], we say that $h$ is $\gamma$-calibrated if for all functions $w:[0,1]\to[0,1]$,
$\Bigl\lvert\operatorname*{\mathbb{E}}_{x\sim\mathcal{D}}[w(h(x))(g(x)-h(x)]\Bigr\rvert\leq\gamma.$
The following known result requires only slight modifications to the proof of Lemma 2.4. We will state it in terms of the class $\mathcal{F}_{(s_{1},s_{2})}$ of functions $h$ with the following property: there exist functions $f_{1},\ldots,f_{k}\in\mathcal{F}$ with $k\leq s_{1}$ and a Boolean circuit of size at most $s_{2}$ that computes the output of $h(x)$ in binary given only $f_{1}(x),\ldots,f_{k}(x)$ as input.
Lemma 2.7.
For all $\mathcal{D}\in\Delta(\mathcal{X})$, $\mathcal{F}\subseteq\{\mathcal{X}\to[-1,1]\}$, $g:\mathcal{X}\to[0,1]$, and $0<\gamma\leq\varepsilon\leq 1$, there exists an $(\mathcal{F},\varepsilon)$-regular and $\gamma$-calibrated simulator in $\mathcal{F}_{(O(1/\varepsilon^{2}),\,\tilde{O}(1/\gamma^{3}))}$.
We have chosen to work with a fairly strong definition of approximate calibration. Various relaxations exist in the literature that would lead to an even milder dependence on $\gamma$ in the complexity of $h$ [GKSZ22]. More stringent than calibration is multicalibration [HKRR18, KNRW18], which is essentially a per-level-set regularity requirement on the simulator. We say that $h$ is $(\mathcal{F},\varepsilon)$-multicalibrated if
[…]
(Several variations of the definition of multicalibration exist in the literature, and the one we have stated is often referred to as swap multicalibration or strict multicalibration.) To emphasize the relationship between $(\mathcal{F},\varepsilon)$-regularity and $(\mathcal{F},\varepsilon)$-multicalibration, we will sometimes follow the now-standard convention of referring to the former as $(\mathcal{F},\varepsilon)$-multiaccuracy, where “accuracy” is short for accuracy in expectation.
One can check that $(\mathcal{F},\varepsilon)$-multicalibration implies both $(\mathcal{F},O(\varepsilon))$-multiaccuracy and, if $\mathcal{F}$ contains the constant 0 and $1$ functions, $O(\varepsilon)$-calibration. Moreover, $(\mathcal{F},\varepsilon)$-multicalibration can be achieved at the cost of a moderate increase in the complexity of $h$ relative to $\mathcal{F}$:
Lemma 2.8 ([HKRR18]).
For all $\mathcal{D}\in\Delta(\mathcal{X})$, $\mathcal{F}\subseteq\{\mathcal{X}\to[-1,1]\}$, $g:\mathcal{X}\to[0,1]$, and $0<\varepsilon\leq 1$, there exists an $(\mathcal{F},\varepsilon)$-multicalibrated simulator in $\mathcal{F}_{(O(1/\varepsilon^{4}),\,\tilde{O}(1/\varepsilon^{4}))}$.
We emphasize that Lemma 2.8 will not be used directly in any of our proofs, but it is nevertheless a useful point of reference. We also remark all of these concepts have been studied in the so-called multiclass case, corresponding to $g:\mathcal{X}\to\Delta([k])$. In this case, which we do not study in this paper, the complexity of known multicalibration constructions are dramatically worse, degrading exponentially with $k$, motivating the study of whether more lightweight regularity notions suffice for downstream applications.
Graph Regularity Iteration
Given a simple graph $G=(V,E)$, let $g=\bm{1}_{E}:V\times V\to\{0,1\}$ denote its edge indicator function, and let $\mathcal{F}=\{\bm{1}_{S\times T}:S,T\subseteq V\}$ contain the indicator functions for all rectangles $S\times T\subseteq V\times V$, sometimes called cuts. Famous and well-studied notions of graph regularity, such as Szemerédi regularity [Sze75] and Frieze-Kannan regularity [FK96, FK99], are closely related to multicalibration and multiaccuracy with respect to $\mathcal{F}$, respectively [TTV09, Skó17, DLLT23]. This connection suggests that techniques developed in the graph regularity literature may have counterparts in our complexity-theoretic setting.
We confirm this intuition in Appendix A, where we adapt the technique of iteration, which constructs a sequence of regular partitions $\mathcal{P}_{1},\mathcal{P}_{2},\ldots$ of the same graph $G$. In this method, each $\mathcal{P}_{i+1}$ is a refinement of $\mathcal{P}_{i}$, and each $\mathcal{P}_{i+1}$ is extremely regular relative to the complexity of $\mathcal{P}_{i}$ (more formally, $\mathcal{P}_{i+1}$ achieves regularity with an error parameter $\varepsilon_{\lvert\mathcal{P}_{i}\rvert}$ that decays with the number of parts of $\mathcal{P}_{i}$). With some care, one can ensure that there exist consecutive partitions $\mathcal{P}=\mathcal{P}_{i^{\star}}$ and $\mathcal{Q}=\mathcal{P}_{i^{\star}+1}$, where $i^{\star}\leq O(1/\varepsilon^{2})$, such that the $\mathcal{P}$ and $\mathcal{Q}$ are distance at most $\varepsilon$ in an appropriately defined metric. Depending on the base regularity notion chosen to instantiate this technique, the resulting pair $(\mathcal{P},\mathcal{Q})$ may have substantially stronger regularity properties than $\mathcal{P}$ alone, allowing one to derive, for example, Szemerédi regularity from the weaker notion of Frieze-Kannan regularity, as well as improved bounds for various combinatorial applications. For more detail about this technique in the graph setting, we refer the reader to [Zha23], as well as [RS10, Fox11, CF12].
Graph Regularity Templates
In Section 1.2.1, we defined regular partitions, regularity templates, and regular-reducibility. In order to prove our results regarding node-level private graph property testing, we will need a few more technical definitions and results regarding these concepts. The first lemma that we need measures the number of edge edits required to improve regularity by a desired amount $\delta$.
Lemma 2.9 ([AFNS06]).
Let $R=(\gamma,k,\eta,\bar{R})$ be a regularity template and let $(V_{1},\ldots,V_{k})$ be an equipartition of a graph $G$ such that $\lvert d(V_{i},V_{j})-\eta_{ij}\rvert\leq\delta$ for all pairs $(i,j)\in[k]^{2}$ and $(V_{i},V_{j})$ is $(\gamma+\delta)$-regular for all pairs $(i,j)\notin\bar{R}$. Then $G$ is $O(\delta/\gamma^{2})$-close to satisfying $R$.
The next definition that we need is a notion of similarity between two vertex equipartitions (not necessarily of the same underlying graph).
Definition 2.10 ([AFNS06]).
An equipartition $\mathcal{U}=\{U_{1},\ldots,U_{k}\}$ of a graph $G$ is $\delta$-similar to a $\gamma$-regular equipartition $\mathcal{V}=\{V_{1},\ldots,V_{k}\}$ of a graph $H$ if $\lvert d_{G}(U_{i},U_{j})-d_{H}(V_{i},V_{j})\rvert\leq\delta$ for all pairs $(i,j)\in[k]^{2}$ and if $(U_{i},U_{j})$ is $(\gamma+\delta)$-regular whenever $(V_{i},V_{j})$ is $\gamma$-regular.
The next lemma states that one can identify regular partitions of a graph by analyzing a constant-size subsample of the vertices. This lemma is intuitive but certainly not trivial, since it is far from obvious that so few samples suffice for this purpose.
Lemma 2.11 ([AFNS06]).
For all $r\in\mathbb{N}$ and $\delta>0$, there exists $q=O_{r,\delta}(1)$ such that if $\gamma\geq\delta$ and $k\leq r$, then the following property holds with probability at least $2/3$ over a uniformly random sample $Q$ of $q$ vertices from a graph $G$: for every size-$k$ $\gamma$-regular equipartition $\mathcal{V}$ of $G$, there exists a size-$k$ equipartition $\mathcal{U}$ of $G[Q]$ that is $\delta$-similar to $\mathcal{V}$, and for every size-$k$ $\gamma$-regular equipartition $\mathcal{U}$ of $G[Q]$, there exists a size-$k$ equipartition $\mathcal{V}$ of $G$ that is $\delta$-similar to $\mathcal{U}$.
The next lemma asserts that any graph property that is testable from from a constant number of edge queries is also testable by inspecting the subgraph induced by a constant number of randomly sampled vertices. The intuition behind the lemma is that a graph property is by definition invariant under vertex relabeling, so it should not matter which part of the graph we inspect.
Lemma 2.12 ([AFKS99, GT01]).
If property $\mathcal{P}$ of $n$-vertex graphs is testable for all $\varepsilon>0$ with proximity $\varepsilon$ and $q/2=O_{\varepsilon}(1)$ edge queries, then $\mathcal{P}$ is testable for all $\varepsilon>0$ with a tester that first selects a uniformly random subset $Q$ of $q$ vertices from $G$, then queries all the pairs in $Q$, and finally outputs Accept or Reject in a deterministic manner that depends only on the isomorphism class of the induced subgraph $G[Q]$.
The final result we need is the equivalence between tolerant and non-tolerant property testing in the dense graph model.
Lemma 2.13 ([FN05]).
If for all $\varepsilon>0$, a graph property $\mathcal{P}$ is testable with proximity $\varepsilon$ using $O_{\varepsilon}(1)$ edge queries, then for all $\varepsilon>0$ and $\delta\geq 0$, the graph property $\mathcal{P}_{\delta}$ is testable with proximity $\varepsilon$ using $O_{\varepsilon,\delta}(1)$ edge queries (i.e. $\mathcal{P}$ is tolerantly testable with proximity parameters $\delta$ and $\delta+\varepsilon$).
Differential Privacy
For our results on private graph regularity, we will need the definition of DP from [DMNS06], specialized to node-adjacent graphs, as in [BBDS13, CZ13, KNRS13], the utility guarantees for the classic exponential mechanism [MT07], and the powerful technique of privacy amplification by subsampling (see [BBG18, Ste22] for an overview of this technique).
Definition 2.14 (Node-Level Differential Privacy [BBDS13, CZ13, KNRS13]).
Two graphs $G,G^{\prime}$ are node-adjacent if $G^{\prime}$ can be obtained by selecting a single vertex $v$ and arbitrarily adding or removing edges from $G$ that are incident to $v$. A randomized algorithm $\mathcal{A}$ that takes as input a graph and outputs an element of a set $\mathcal{S}$ satisfies node-level $\alpha$-differential privacy ($\alpha$-DP) if for all node-adjacent graphs $G,G^{\prime}$ and all possible outputs $s\in\mathcal{S}$,
$\Pr[\mathcal{A}(G)=s]\leq e^{\alpha}\Pr[\mathcal{A}(G^{\prime})=s],$
where the probabilities are computed over only the randomness internal to the algorithm $\mathcal{A}$.
Lemma 2.15 (Exponential Mechanism [MT07]).
Let $\ell_{1},\ell_{2}$ be two real-valued functions on a finite domain $\mathcal{R}$ that differ pointwise by at most $\Delta$, and let $\mathcal{D}_{i}$ be the distribution over $\mathcal{R}$ proportional to $\exp(-(\alpha/2\Delta)\cdot\ell_{i}(R))$. Then, $\mathcal{D}_{1}(R)\leq e^{\alpha}\cdot\mathcal{D}_{2}(R)$ for each $R\in\mathcal{R}$, and for all $\beta>0$,
$\Pr_{R\sim\mathcal{D}_{1}}\Bigl[\ell_{1}(R)\leq\min_{R^{*}\in\mathcal{R}}\ell_{1}(R^{*})+\frac{2\Delta}{\alpha}\log\Bigl(\frac{\lvert\mathcal{R}\rvert}{\beta}\Bigr)\Bigr]\geq 1-\beta.$
Lemma 2.16 (Privacy Amplification by Subsampling, Variant of Theorem 29 of [Ste22]).
Let $Q\subseteq[n]$ be a random subset. Assume that, for all $i\in[n]$, we can define a random subset $Q_{-i}\subseteq[n]\setminus\{i\}$ distributed jointly with $Q$ such that the following two conditions hold for all $i\in[n]$:
  • For all $n$-vertex graphs $G$, the induced subgraphs $G[Q]$ and $G[Q_{-i}]$ are node-adjacent with probability $1$ over the randomness of $Q$ and $Q_{-i}$,
  • The marginal distribution of $Q_{-i}$ conditioned on $i\in Q$ is equal to the marginal distribution of $Q$ conditioned on $i\notin Q$.
Let $M$ be an $\alpha$-DP mechanism that takes as input an $n$-vertex graph. Define the mechanism $M^{Q}(G)=M(G[Q])$. Let $p=\max_{i\in[n]}\Pr_{Q}[i\in Q]$. Then $M^{Q}$ is $\alpha^{\prime}$-DP for $\alpha^{\prime}=\log(1+p(e^{\alpha}-1))$.
Whenever we use Lemma 2.16, we will let $Q$ be a uniformly random subset of $[n]$ of size exactly $q$, for some desired size parameter $q\in[n]$. We will let $Q_{-i}$ be the result of replacing the element $i$ (if present in $Q$) with a uniformly random element in $n]\setminus Q$, as described in [Ste22].